# LoXR: Performance Evaluation of Locally Executing LLMs on XR Devices

### v1

@@ 4.1 Models Quality Analysis
Due to the storage and computational limitations of edge devices, deploying LLM locally often requires a trade-off between model size and performance. To put our performance results in perspective, we first conducted an empirical analysis of the models we have used in this study. This section offers insights into the models evaluated in this work, presenting a comparative analysis. While this section is not directly related to device performance, we believe providing detailed information about the models is essential for understanding their capabilities and limitations. We select six benchmarks to evaluate each model and calculate different metrics for each run:
  • Hellaswag [24]: commonsense reasoning tasks, focusing on selecting the most plausible ending to a given situation from multiple options.
  • mmlu [25]: evaluating a model’s knowledge and reasoning across 57 different subjects, ranging from elementary-level topics to advanced professional knowledge.
  • ARC [26]: evaluating model’s ability to answer complex, grade-school science questions that require reasoning and problem-solving beyond simple fact retrieval.
  • Truthful-qa [27]: evaluate how accurately AI models generate truthful answers to questions, especially in cases where common misconceptions or false information could lead to incorrect responses.
  • winogrande [28]: large-scale benchmark for commonsense reasoning, specifically designed to test a model’s ability to resolve ambiguous pronouns in sentences.
  • WikiText-2 [29]: containing over 100 million tokens, sourced from verified Wikipedia articles, designed to support research in language modeling and text generation tasks.
@@
$\text{Perplexity}(P)=e^{-\frac{1}{N}\sum_{i=1}^{N}\log P(w_{i})}$ (4)
@@ 4.2 Evaluating Performance Consistency
Some devices may experience performance degradation over time. In these experiments, we evaluate how processing speed fluctuates across multiple runs. For each model $m_{i}$ and device $d_{j}$, we perform $N$ runs and compute the mean $(\mu_{ij})$ of the speed $(S_{c})$, standard deviation $\sigma_{ij}$, and coefficient of variation ($CV_{ij}$). We also record the error count, as well as the maximum and minimum speeds observed during each experiment.
In our experiments, we set $N=20$, i.e., each model-device pair is tested 20 times to evaluate performance variance. The mean speed across $N$ runs is calculated as:
$\mu=\frac{\sum_{i=1}^{N}S_{c}}{N},$ (5)
where $S_{c}$ represents the 5-Run-Mean speed measured during each run, calculated as in Eq. 3. In other words, each $S_{c}$ itself is the mean of 5 runs, ensuring that the measured value is not too affected by errors or hidden factors.
To let the devices cool down and conduct each run in comparable circumstances, we leave each device idle for two minutes between each run of the same model. We find this to be adequate, but just to be safer, between different models, we leave each device idle for ten minutes. We noted that the Meta Quest is not very prone to heating up. Heating was more noticeable on the Vivo X100s Pro, so it is always placed on a metal plate for heat dissipation. This is not representative of actual use, but this device is included in our evaluation as a baseline, not as an example of an actual XR headset. Magic Leap features a cooling fan, as does Vision Pro.
For each test, we conducted both PP and TG evaluations with the string length set to 64, i.e., $\text{PP}=64$ and $\text{TG}=64$. For each model-device pair, we calculated the standard deviation, mean, and CV. The results are presented in Section 5.2.
@@ 4.3 Processing Speed and String Length
  • PP: The PP values were set to 64, 128, 256, 512, and 1024, while TG was fixed at 0, with all other parameters set to their default values.
  • TG: The TG values were set to 64, 128, 256, 512, and 1024, while PP was fixed at 0, with all other parameters set to their default values.
The results are shown in Section 5.3.
@@ 4.4 Parallelization with Thread Count and Batch Size
  • BT: We set PP to 64, TG to 0, and varied the batch sizes between 128, 256, 512, and 1024, with all other parameters kept at their default values.
  • TT: We set PP to 64 and TG to 0, while varying the thread counts between 1, 2, 4, 8, 16, and 32. All other parameters were kept at their default values.
The results are shown in Section 5.4.
@@ 4.7 Pareto Optimality
where $o$ is the objective index, $n$ is the number of metrics used in the objective category, $w_{i}$ defines the weight for each metric, and $\min$ and $\max$ calculate the minimum and maximum values across the entire set of results for metric $m_{i}$.
In practice, some metrics, such as perplexity, are better when lower, which is the opposite of the direction in Eq. 6. Therefore, we take the reciprocal of each perplexity value.
Figure 4: Consistency results of the Apple Vision Pro on GPU. Left: PP results. Right: TG results. Each model was tested 20 times, and the results are plotted in sorted order.
Table II: Performance consistency results for the four devices: Magic Leap 2 (ML2), Meta Quest 3 (MQ3), Vivo X100 Pro (Vivo), and Apple Vision Pro (VPro for CPU and VPro* for GPU). Processing speed was calculated from 20 runs for each model-device pair. The results are reported in terms of the mean ($\mu$) speed (t/s) of 20 runs, its standard deviation ($\sigma$), coefficient of variation (CV %), and the range of values: $[min,max]$.
D Metric m2 m3 m4 m5 m6 m7 m8 m9 m10 m11 m12 m13 m14 m15 m16 m17
PP Results ML2 $\mu$±$\pm\sigma$ $17.01\pm 4.67$ $13.99\pm 4.57$ $10.38\pm 2.74$ $20.82\pm 0.87$ $16.81\pm 0.07$ $13.16\pm 0.47$ $14.22\pm 0.31$ $12.63\pm 0.38$ $13.12\pm 0.29$ $11.45\pm 0.18$ $7.92\pm 0.05$ $7.37\pm 0.19$ $2.20\pm 0.03$ $3.14\pm 0.08$ $1.15\pm 0.05$ $1.52\pm 0.09$
Range $[10.25,22.24]$ $[7.47,18.75]$ $[6.27,13.32]$ $[18.28,21.36]$ $[16.68,16.95]$ $[12.04,13.51]$ $[12.98,14.36]$ $[11.73,12.87]$ $[12.00,13.27]$ $[10.93,11.62]$ $[7.73,7.99]$ $[6.73,7.54]$ $[2.14,2.23]$ $[2.98,3.27]$ $[1.10,1.33]$ $[1.36,1.72]$
CV (%) $27.46$ $32.66$ $26.41$ $4.16$ $0.43$ $3.56$ $2.19$ $2.98$ $2.22$ $1.57$ $0.66$ $2.54$ $1.35$ $2.41$ $4.70$ $5.99$
MQ3 $\mu$±$\pm\sigma$ $21.00\pm 2.16$ $25.61\pm 0.10$ $21.33\pm 1.47$ $16.33\pm 1.84$ $10.99\pm 1.65$ $11.37\pm 1.80$ $13.74\pm 1.72$ $10.24\pm 1.34$ $11.07\pm 0.61$ $8.17\pm 2.19$ $7.12\pm 0.04$ $6.09\pm 0.32$ $2.65\pm 0.21$ $2.37\pm 0.12$ $1.57\pm 0.06$ $2.04\pm 0.09$
Range $[17.42,23.59]$ $[25.03,25.77]$ $[14.79,23.12]$ $[13.66,18.72]$ $[7.62,13.89]$ $[7.91,13.86]$ $[10.88,16.74]$ $[7.79,11.87]$ $[7.19,11.28]$ $[4.31,13.05]$ $[5.96,7.17]$ $[5.64,6.61]$ $[2.10,2.92]$ $[2.12,2.57]$ $[1.44,1.68]$ $[1.85,2.34]$
CV (%) $10.27$ $0.37$ $6.89$ $11.24$ $14.99$ $15.80$ $12.51$ $13.06$ $5.55$ $26.81$ $0.56$ $5.23$ $7.84$ $5.24$ $3.93$ $4.23$
Vivo $\mu$±$\pm\sigma$ $15.70\pm 2.88$ $15.90\pm 2.76$ $14.98\pm 2.98$ $13.77\pm 2.13$ $8.66\pm 0.80$ $8.27\pm 1.08$ $10.35\pm 1.10$ $8.48\pm 1.08$ $8.82\pm 1.61$ $11.14\pm 1.21$ $4.47\pm 0.19$ $4.35\pm 0.23$ $2.44\pm 0.07$ $4.28\pm 0.44$ $2.61\pm 0.12$ $5.89\pm 0.47$
Range $[11.16,25.55]$ $[9.39,17.81]$ $[10.52,24.36]$ $[10.65,15.55]$ $[7.32,9.34]$ $[5.88,9.53]$ $[8.27,11.05]$ $[7.19,11.72]$ $[5.97,11.60]$ $[9.13,12.04]$ $[3.98,4.92]$ $[3.80,4.86]$ $[2.27,2.52]$ $[2.61,4.80]$ $[2.23,2.76]$ $[5.17,6.59]$
CV (%) $18.33$ $17.38$ $19.88$ $15.50$ $9.26$ $13.07$ $10.61$ $12.75$ $18.31$ $10.84$ $4.24$ $5.39$ $3.00$ $10.30$ $4.55$ $7.92$
VPro $\mu$±$\pm\sigma$ $31.03\pm 2.20$ $41.22\pm 3.04$ $33.11\pm 1.35$ $32.96\pm 1.21$ $20.62\pm 0.56$ $19.34\pm 0.64$ $23.14\pm 1.09$ $19.52\pm 0.73$ $19.56\pm 0.60$ $29.22\pm 1.22$ $9.15\pm 0.40$ $9.06\pm 0.28$ $5.52\pm 0.21$ $7.44\pm 0.40$ $5.01\pm 0.26$ $14.05\pm 0.56$
Range $[29,34.98]$ $[39.44,47.03]$ $[32.98,34.79]$ $[32.15,35.93]$ $[20.19,21.63]$ $[19.09,20.43]$ $[22.65,24.92]$ $[19.04,20.98]$ $[19.15,20.74]$ $[28.58,31.25]$ $[8.77,9.67]$ $[8.88,9.77]$ $[5.4,5.93]$ $[7.21,8.13]$ $[4.83,5.47]$ $[13.63,15.02]$
CV (%) $7.08$ $7.37$ $4.07$ $3.68$ $2.72$ $3.31$ $4.69$ $3.72$ $3.05$ $4.17$ $4.41$ $3.12$ $3.73$ $5.38$ $5.16$ $3.99$
VPro* $\mu$±$\pm\sigma$ $378\pm 3.7$ $432\pm 3.7$ $382\pm 4.8$ $374\pm 4.6$ $241\pm 1.5$ $230\pm 2.3$ $237\pm 2.0$ $218\pm 5.8$ $229\pm 4.3$ $264\pm 3.4$ $136\pm 9.9$ $128\pm 12.5$ $135\pm 2.3$ $134\pm 2.3$ $133\pm 1.4$ $130\pm 6.6$
Range $[369,384]$ $[424,439]$ $[367,389]$ $[363,380]$ $[237,243]$ $[225,234]$ $[232,240]$ $[198,223]$ $[221,237]$ $[255,268]$ $[109,145]$ $[100,143]$ $[130,138]$ $[129,136]$ $[129,134]$ $[109,135]$
CV (%) $1.0$ $0.9$ $1.2$ $1.2$ $0.6$ $1.0$ $0.8$ $2.6$ $1.9$ $1.3$ $7.3$ $9.8$ $1.7$ $1.7$ $1.1$ $5.1$
TG Results ML2 $\mu$±$\pm\sigma$ $9.14\pm 1.92$ $8.22\pm 1.92$ $6.53\pm 1.77$ $8.43\pm 0.07$ $11.69\pm 0.04$ $8.52\pm 0.02$ $7.88\pm 0.01$ $6.92\pm 0.01$ $6.25\pm 0.10$ $5.05\pm 0.02$ $5.90\pm 0.03$ $5.59\pm 0.14$ $2.05\pm 0.01$ $2.84\pm 0.06$ $0.24\pm 0.01$ $0.21\pm 0.01$
Range $[6.92,11.25]$ $[5.69,10.27]$ $[4.41,8.39]$ $[8.33,8.55]$ $[11.61,11.74]$ $[8.48,8.55]$ $[7.85,7.91]$ $[6.89,6.95]$ $[5.95,6.31]$ $[5.01,5.08]$ $[5.83,5.94]$ $[5.13,5.73]$ $[2.04,2.06]$ $[2.77,2.97]$ $[0.22,0.27]$ $[0.19,0.23]$
CV (%) $20.95$ $23.36$ $27.10$ $0.77$ $0.32$ $0.23$ $0.18$ $0.21$ $1.62$ $0.39$ $0.48$ $2.49$ $0.32$ $2.24$ $5.32$ $5.20$
MQ3 $\mu$±$\pm\sigma$ $12.37\pm 1.45$ $14.19\pm 0.30$ $11.58\pm 0.98$ $9.47\pm 1.29$ $9.00\pm 1.12$ $8.57\pm 0.70$ $9.30\pm 0.51$ $7.61\pm 0.65$ $8.30\pm 0.26$ $5.47\pm 0.10$ $5.62\pm 0.23$ $5.01\pm 0.24$ $2.95\pm 0.21$ $3.06\pm 0.05$ $1.67\pm 0.15$ $3.38\pm 0.09$
Range $[9.11,13.89]$ $[13.52,14.67]$ $[7.77,12.80]$ $[7.64,10.87]$ $[7.38,10.39]$ $[6.85,9.55]$ $[7.66,9.86]$ $[6.13,8.24]$ $[7.40,8.53]$ $[5.19,5.61]$ $[5.26,6.17]$ $[4.61,5.57]$ $[2.23,3.40]$ $[2.96,3.13]$ $[1.40,1.84]$ $[3.11,3.53]$
CV (%) $11.73$ $2.11$ $8.43$ $13.65$ $12.50$ $8.17$ $5.52$ $8.51$ $3.18$ $1.87$ $4.09$ $4.83$ $7.27$ $1.68$ $8.93$ $2.80$
Vivo $\mu$±$\pm\sigma$ $9.39\pm 0.98$ $9.85\pm 0.42$ $9.44\pm 1.22$ $8.66\pm 0.30$ $6.91\pm 0.08$ $6.83\pm 0.22$ $8.21\pm 0.12$ $6.84\pm 0.61$ $6.61\pm 0.71$ $7.51\pm 0.11$ $3.70\pm 0.05$ $3.59\pm 0.11$ $2.09\pm 0.01$ $3.71\pm 0.04$ $2.27\pm 0.02$ $4.72\pm 0.17$
Range $[8.54,13.35]$ $[9.17,10.67]$ $[7.58,13.86]$ $[7.97,9.26]$ $[6.75,7.00]$ $[6.39,7.21]$ $[7.52,8.35]$ $[4.37,7.11]$ $[5.70,7.68]$ $[7.25,7.64]$ $[3.60,3.77]$ $[3.41,3.71]$ $[2.06,2.11]$ $[3.65,3.79]$ $[2.23,2.31]$ $[4.48,5.00]$
CV (%) $10.46$ $4.22$ $12.91$ $3.48$ $1.13$ $3.25$ $1.48$ $8.92$ $10.75$ $1.41$ $1.27$ $3.04$ $0.68$ $1.15$ $1.06$ $3.61$
VPro $\mu$±$\pm\sigma$ $17.90\pm 1.00$ $21.15\pm 0.82$ $18.92\pm 0.74$ $17.71\pm 0.50$ $15.72\pm 0.37$ $14.76\pm 0.26$ $16.63\pm 0.49$ $13.93\pm 0.50$ $14.07\pm 0.19$ $12.03\pm 0.21$ $7.20\pm 0.21$ $7.22\pm 0.17$ $4.53\pm 0.17$ $6.24\pm 0.27$ $4.17\pm 0.18$ $10.22\pm 0.30$
Range $[16.59,19.72]$ $[19.4,22.49]$ $[17.31,19.8]$ $[16.89,18.77]$ $[14.88,16.36]$ $[14.32,15.28]$ $[15.77,17.56]$ $[12.89,14.99]$ $[13.71,14.49]$ $[11.26,12.28]$ $[6.74,7.52]$ $[6.98,7.7]$ $[4.11,4.93]$ $[5.52,6.72]$ $[3.86,4.52]$ $[9.82,10.93]$
CV (%) $5.58$ $3.86$ $3.89$ $2.84$ $2.35$ $1.73$ $2.97$ $3.61$ $1.35$ $1.77$ $2.95$ $2.41$ $3.84$ $4.35$ $4.31$ $2.91$
VPro* $\mu$±$\pm\sigma$ $23.1\pm 0.1$ $27.5\pm 0.1$ $23.5\pm 0.1$ $22.6\pm 0.1$ $23.3\pm 0.1$ $19.6\pm 0.8$ $19.9\pm 0.1$ $18.0\pm 0.4$ $17.7\pm 0.2$ $14.6\pm 0.2$ $13.7\pm 0.8$ $12.5\pm 1.0$ $14.5\pm 0.3$ $15.1\pm 0.2$ $15.2\pm 0.2$ $14.7\pm 0.4$
Range $[23,23.3]$ $[27.1,27.7]$ $[23.2,23.8]$ $[22.3,22.7]$ $[23.0,23.5]$ $[19.2,22.8]$ $[19.8,20.2]$ $[16.8,18.5]$ $[17.2,17.9]$ $[14.3,14.9]$ $[11.4,14.6]$ $[10.2,13.9]$ $[13.9,14.9]$ $[14.6,15.4]$ $[14.9,15.4]$ $[13.5,15.1]$
CV(%) $0.4$ $0.5$ $0.5$ $0.3$ $0.6$ $4.3$ $0.6$ $2.1$ $1.0$ $1.1$ $5.8$ $8.3$ $1.9$ $1.4$ $1.1$ $2.5$
@@ 5.2 Performance Consistency Results
Figures 3 and 4 illustrate the consistency of performance across all model-device pairs, while Table II provides the corresponding quantitative results. Across all 20 runs, the Apple Vision Pro demonstrates the most consistent results compared to other devices, with the corresponding speed also being the fastest. Notably, the GPU results for the Apple Vision Pro exhibit greater stability and smaller variance. Among the remaining devices, the Magic Leap 2 also shows a reasonable degree of stability; however, for the first three models, its variance is surprisingly high, reaching 27%, 32%, and 26% for $m_{2}$, $m_{3}$, and $m_{3}$ under the PP setup, and 20%, 23%, and 27% under the TG setup, respectively. In contrast, the Meta Quest 3 and Vivo X100 Pro exhibit relatively poor performance both in terms of speed and variance. Between these two devices, the Meta Quest 3 achieves higher speeds than the Vivo X100 Pro, but there is no clear difference in variance. For some models, one device shows lower variance, while for others, the variance is higher.
Furthermore, as shown in Table III performance varies significantly for $m_{1}$ over Magic Leap 2 and Meta Quest 3, but it is generally consistent across the other two devices. These findings suggest that, provided model-device pairs are carefully selected, LLM performance can be considered reliable, strengthening the case for on-device LLM use in XR environments.
Table III: Performance consistency results for $m_{1}$ across the four devices. For the results of other models and description of metrics used, see Table II.
Test Metric ML2 Vivoo Meta Q3 VisPro VisPro*
PP $\mu\pm\sigma$ 51.68 $\pm$ 4.22 62.68 $\pm$ 16.64 39.32 $\pm$ 15.66 292.48 $\pm$ 12.87 1603.59 $\pm$ 56.12
Range [43.62, 54.59] [21.63, 72.04] [24.25, 72.59] [288.5, 312.1] [1516.9, 1696.1]
CV (%) 8.17 26.55 39.82 4.40 3.50
TG $\mu\pm\sigma$ 20.26 $\pm$ 0.05 22.16 $\pm$ 0.90 16.76 $\pm$ 3.76 41.94 $\pm$ 1.77 42.94 $\pm$ 0.80
Range [20.17, 20.34] [20.54, 23.41] [11.61, 22.53] [38.90, 46.37] [41.76, 44.35]
CV (%) 0.23 4.05 22.44 4.21 1.87
Figure 5: Processing speed of the four devices in the PP test (top) and TG test (bottom) with varying string lengths: 64, 128, 256, 512, and 1024. The x-axis represents the model, while the y-axis represents the processing speed in t/s.
Figure 6: Processing speed of Apple Vision Pro (GPU) in the PP test (left) and TG test (rigth). The x-axis represents the model, while the y-axis represents the processing speed in t/s.
Figure 7: Error counts for Meta Quest 3 (left) and Apple Vision Pro (right) in PP (top) and TG (bottom). Magic Leap 2, with zero errors, and Vivo X100 Pro, with a total of four errors, one in $m_{1}$ (PP-64), two in $m_{10}$ (PP-64, PP-128), and one in $m_{11}$ (TG-1024) are excluded.
@@
Figure 9: Thread Test (TT) results: The X-axis represents the thread count, while the Y-axis shows the speed in tokens per second. Threads 4 and 8 deliver the fastest results. Note: Apple Vision Pro (CPU) fails for thread counts of 16 and 32.
5.3 Results of Processing Speed and String Length
Figs. 5 and 6 present the processing speed results of PP and TG with varying string lengths (prompt length and token sets). Both PP and TG were tested with values of 64, 128, 256, 512, and 1024. The processing speed varies significantly across string lengths, devices, and models.
@@ 5.3 Results of Processing Speed and String Length
Error Counts: Figure 7 shows the errors counts during the PP and TG tests. Errors mostly occurred in AVP (CPU) and MQ3, particularly at longer prompt lengths of 512 and 1024 tokens. There are a few reasons for this. First, longer strings process more slowly, leading to abrupt changes and occasional failures. Second, memory constraints or inefficiencies in sustained generation contribute to these errors. With AVP, the general user experience is also suboptimal, as the device must be actively mounted on the head, which can cause errors if not handled carefully. Overall, MQ3 records the highest count of errors, with frequent retries affecting its reliability, whereas ML2 remains highly stable, leading to minimal inconsistencies (zero errors). As expected, AVP (GPU) also reported zero errors.

### v2

@@ 4.1 Models Quality Analysis
Due to the storage and computational limitations of edge devices, deploying LLM locally often requires a trade-off between model size and performance. To put our performance results in perspective, we first conducted an empirical analysis of the models we have used in this study. This section offers insights into the models evaluated in this work, presenting a comparative analysis. While this section is not directly related to device performance, we believe providing detailed information about the models is essential for understanding their capabilities and limitations. We evaluate each model using six benchmarks, assessing various reasoning and knowledge capabilities: Hellaswag [48] (commonsense reasoning), MMLU [49] (multi-subject knowledge), ARC [50] (complex science questions), TruthfulQA [51] (truthfulness in responses), Winogrande [52] (pronoun resolution), and WikiText-2 [53] (language modeling and text generation).
@@
$\text{Perplexity}(P)=e^{-\frac{1}{N}\sum_{i=1}^{N}\log P(w_{i})},$ (2)
@@ 4.2 Evaluating Performance Consistency
Some devices may experience performance degradation over time. In these experiments, we evaluate how processing speed fluctuates across multiple runs. For each model $m_{i}$ and device $d_{j}$, we perform $N$ runs and compute the mean $(\mu_{ij})$ of the speed $(S_{c})$, standard deviation $\sigma_{ij}$, and coefficient of variation ($CV_{ij}$). We also record the error count, as well as the maximum and minimum speeds observed during each experiment. In our experiments, we set $N=20$, i.e., each model-device pair is tested 20 times to evaluate performance variance. We noted the 5-Run-Mean speed $S_{c}$ measured during each run, calculated as in Eq. 1. In other words, each $S_{c}$ itself is the mean of 5 runs, ensuring that the measured value is not too affected by errors or hidden factors.
To let the devices cool down and conduct each run in comparable circumstances, we leave each device idle for two minutes between each run of the same model. We find this to be adequate, but just to be safer, between different models, we leave each device idle for ten minutes. For each test, we conducted both PP and TG evaluations with the string length set to 64, i.e., $\text{PP}=64$ and $\text{TG}=64$. For each model-device pair, we calculated the standard deviation, mean, and CV. The results are presented in Section 5.2.
@@ 4.3 Processing Speed and String Length
  • PP: The PP values were set to 64, 128, 256, 512, and 1024, while TG was fixed at 0, and other parameters at their default values.
  • TG: The TG values were set to 64, 128, 256, 512, and 1024, while PP was set 0, and other parameters were at default values.
Figure 4: Consistency results of the AVP on GPU. Left: PP results. Right: TG results. Each model was tested 20 times, and the results are plotted in sorted order.
@@ 4.4 Parallelization with Thread Count and Batch Size
  • BT: We set PP to 64, TG to 0, and varied the batch size (128, 256, 512, 1024), keeping other parameters at default values.
  • TT: We set PP to 64 and TG to 0, and varied the thread counts (1, 2, 4, 8, 16, 32), keeping other parameters at default values.
@@ 4.7 Pareto Optimality
where $o$ is the objective index, $n$ is the number of metrics used in the objective category, $w_{i}$ defines the weight for each metric, and $\min$ and $\max$ calculate the minimum and maximum values across the entire set of results for metric $m_{i}$. In practice, some metrics, such as perplexity, are better when lower, which is the opposite of the direction in Eq. 3. Therefore, we take the reciprocal of each perplexity value.
Figure 5: Processing speed of the four devices in the PP test (top) and TG test (bottom) with varying string lengths: 64, 128, 256, 512, and 1024. The x-axis represents the model, while the y-axis represents the processing speed in t/s. For $m_{1}$, a different scale is used due to its higher values.
@@ 5.2 Performance Consistency Results
Figs. 3 and 4 illustrate the consistency of performance across all model-device pairs. Across all 20 runs, the AVP demonstrates the most consistent results compared to other devices, with the corresponding speed also being the fastest. Notably, the GPU results for the AVP exhibit greater stability and smaller variance. Among the remaining devices, the Magic Leap 2 also shows a reasonable degree of stability; however, for the first three models, its variance is surprisingly high, reaching 27%, 32%, and 26% for $m_{2}$, $m_{3}$, and $m_{3}$ under the PP setup, and 20%, 23%, and 27% under the TG setup, respectively. In contrast, the Meta Quest 3 and Vivo X100 Pro exhibit relatively poor performance both in terms of speed and variance. Between these two devices, the Meta Quest 3 achieves higher speeds than the Vivo X100 Pro, but there is no clear difference in variance. For some models, one device shows lower variance, while for others, the variance is higher.
5.3 Results of Processing Speed and String Length
Fig. 5 presents the processing speed results of PP and TG with varying string lengths (prompt length and token sets). Both PP and TG were tested with values of 64, 128, 256, 512, and 1024. The processing speed varies significantly across string lengths, devices, and models.
@@ 5.3 Results of Processing Speed and String Length
Figure 7: Thread Test (TT) results: The X-axis represents the thread count, while the Y-axis shows the speed in tokens per second. Threads 4 and 8 deliver the fastest results. Note: AVP (CPU) fails for thread counts of 16 and 32.
Figure 8: Error counts for Meta Quest 3 (left) and AVP (right) in PP (top) and TG (bottom). Magic Leap 2, with zero errors, and Vivo X100 Pro, with a total of four errors, one in $m_{1}$ (PP-64), two in $m_{10}$ (PP-64, PP-128), and one in $m_{11}$ (TG-1024) are excluded.
@@
Error Counts: Fig. 8 shows the errors counts during the PP and TG tests. Errors mostly occurred in AVP (CPU) and MQ3, particularly at longer prompt lengths of 512 and 1024 tokens. There are a few reasons for this. First, longer strings process more slowly, leading to abrupt changes and occasional failures. Second, memory constraints or inefficiencies in sustained generation contribute to these errors. With AVP, the general user experience is also suboptimal, as the device must be actively mounted on the head, which can cause errors if not handled carefully. Overall, MQ3 records the highest count of errors, with frequent retries affecting its reliability, whereas ML2 remains highly stable, leading to minimal inconsistencies (zero errors). As expected, AVP (GPU) also reported zero errors.